# Tackling Graph Oversquashing by Global and Local Non-Dissipativity

### v1

@@ D.1 Graph Transfer Dataset
We built the graph transfer datasets upon (Di Giovanni et al., 2023). In each task, graphs use identical topology, but, differently from the original work, nodes are initialized with random input features sampled from a uniform distribution in the interval $[0,0.5)$. In each graph, we selected a source node and target node and initialized them with labels of value “1” and “0”, respectively. We sampled graphs from three graph distributions, i.e., line, ring, and crossed-ring. Figure 4 shows a visual exemplification of the three types of graphs when the distance between the source and target nodes is 5. Specifically, ring graphs are cycles of size $n$, in which the target and source nodes are placed at a distance of $\lfloor n/2\rfloor$ from each other. Crossed-ring graphs are also cycles of size $n$, but include crosses between intermediate nodes. Even in this case, the distance between source and target nodes remains $\lfloor n/2\rfloor$. Lastly, the line graph contains a path of length $n$ between the source and target node. In our experiments, we consider a regression task, whose aim is to swap source and target node labels while maintaining intermediate nodes unchanged. We use an input dimension of 1, and the distance between source and target nodes is equal to 3, 5, 10, and 50. We generated 1000 graphs for training, 100 for validation, and 100 for testing.
@@ D.2 Graph Property Prediction
In our experiments on graph property prediction, we followed the data generation procedure outlined in (Gravina et al., 2023). Graphs were randomly selected from various graph distributions, including Erdős–R’enyi, Barabasi-Albert, tree, and grid. Each graph contained between 25 and 35 nodes, where each node is assigned with random identifiers as input features sampled from a uniform distribution in the interval $[0,1)$. The target values represented single-source shortest paths, node eccentricity, and graph diameter. The dataset comprised a total of 7040 graphs, with 5120 used for training, 640 for validation, and 1280 for testing.
@@ D.3 Long Range Graph Benchmark
In the Long Range Graph Benchmark section, we considered the “Peptides-func” and “Peptides-struct” datasets (Dwivedi et al., 2022). The graphs correspond to 1D amino acid chains, and they are derived such that the nodes correspond to the heavy (non-hydrogen) atoms of the peptides while the edges represent the bonds between them. Peptides-func is a multi-label graph classification dataset, with a total of 10 classes based on the peptide function, e.g., Antibacterial, Antiviral, cell-cell communication, and others. Peptides-struct is a multi-label graph regression dataset based on the 3D structure of the peptides. Specifically, the task consists of the prediction of the inertia of the molecules using the mass and valence of the atoms, the maximum distance between each atom-pairs, sphericity, and the average distance of all heavy atoms from the plane of best fit. Both Peptides-func and Peptides-struct consist of 15,535 graphs with a total of 2.3 million nodes. We applied stratified splitting to generate balanced train–valid–test dataset splits, using the ratio of 70%–15%–15%, respectively.
@@ Appendix E Experimental Details
Our method is implemented in PyTorch (Paszke et al., 2017) and PyTorch-Geometric (Fey & Lenssen, 2019). All source codes will be released upon publication. Our experimental results were obtained using NVIDIA RTX-3090 and A100 GPUs.
@@ E.1 SWAN versions
In our experiments, we explore four variants of our SWAN method, depending on the implementation of $\Phi$ and $\Psi$, as reported in Table 4. These versions can be grouped into two main groups, which differ in the resulting non-dissipative behavior. In the first group, which contains SWAN and SWAN-learn both $\Phi$ and $\Psi$ lead to purely imaginary eigenvalues, since we use symmetric graph shift operators and antisymmetric weight matrices, thus allowing for a node- and graph-wise non-dissipative behavior. The last group, which includes SWAN-ne and SWAN-learn-ne, can deviate from being globally non-dissipative. Indeed, if the weight matrix, ${\bf V}$, is an arbitrary matrix, then the eigenvalues of the Jacobian matrix of the system, $\mathbf{J}(t)$, are contained in a neighborhood of the imaginary axis with radius $r\leq||\mathbf{V}||$ (Bauer-Fike’s theorem (Bauer & Fike, 1960)). Although this result does not guarantee that the eigenvalues of the Jacobian are imaginary, in practice it crucially limits their position, limiting the dynamics of the system on the graph to show at most moderate amplification or loss of signals through the graph.
@@ E.3 Graph Property Prediction
We employ the same datasets, hyperparameter space, and experimental setting presented in Gravina et al. (2023). Therefore, we perform hyperparameter tuning via grid search, optimizing the Mean Square Error (MSE), training the models using Adam optimizer for a maximum of 1500 epochs, and early stopping with patience of 100 epochs on the validation error. For each model configuration, we perform 4 training runs with different weight initialization and report the average of the results. We report in Appendix E.5 the grid of hyperparameters exploited for this experiment, while in Appendix D more details about the dataset.
@@ E.4 Long-Range Graph Benchmarks
We employ the same datasets and experimental setting presented in Dwivedi et al. (2022). Therefore, we perform hyperparameter tuning via grid search, optimizing the Average Precision (AP) in the Peptide-func task and the Mean Absolute Error (MAE) in the Peptide-struct task, training the models using AdamW optimizer for a maximum of 300 epochs. For each model configuration, we perform 3 training runs with different weight initialization and report the average of the results. Also, we follow the guidelines in (Dwivedi et al., 2022; Gutteridge et al., 2023) and stay within the 500K parameter budget. We report in Appendix E.5 the grid of hyperparameters exploited for this experiment, while in Appendix D more details about the dataset.
@@ E.5 Hyperparameters
Table 5: The grid of hyper-parameters employed during model selection for the graph transfer tasks (Transfer), graph property prediction tasks (GraphProp), and Long Range Graph Benchmark (LRGB). We observe that, in each graph transfer task, we use a number of layers that is equal to the distance between the source and target nodes.
@@ Appendix F Complexity and Runtimes
Complexity Analysis.
Our SWAN architecture remains within the computational complexity of MPNNs (e.g., Morris et al. (2019); Xu et al. (2019)) and other DE-GNNs such as GRAND (Chamberlain et al., 2021b). Specifically, each SWAN layer is linear in the number of nodes $|\cal{V}|$ and edges $|\cal{E}|$, therefore it has a time complexity of $\mathcal{O}(|\cal{V}|+|\cal{E}|)$. Assuming we compute $L$ steps of the ODE defined by SWAN, the overall complexity is $\mathcal{O}\left(L\cdot(|\cal{V}|+|\cal{E}|)\right)$.
Runtimes. We measure the training and inference (i.e., the epoch-time and test-set time) runtime of SWAN on the Peptides-struct dataset and compare it with other baselines. Our runtimes show that SWAN is able to obtain high performance, while retaining a linear complexity runtimes that is parallel to other MPNNs and DE-GNNs. Our measurements are reported in Table 6, and presents the training and inference runtimes in seconds, as well as the obtained metric for that task, the mean-absolute-error (MAE), for reference. The runtimes were measured on an NVIDIA RTX-3090 GPU with 24GB of memory. In all time measurements, we use a batch size of 64, 128 feature channels, and 5 layers. For reference, we also provide the reported downstream task performance of each method.
Table 6: Measured training and inference runtimes in seconds, and the obtain MAE of SWAN and other baselines on the Peptides-func dataset.
@@ Appendix G Additional Comparisons
In Table 2 we are interested in directly comparing the performance of SWAN and its non-dissipative properties with other MPNNs and DE-GNNs, as well as the common approach of using graph transformer to address long-range interaction modeling. We now provide additional comparisons with other methods that utilize multi-hop information, and are therefore more computationally expensive than our SWAN, while also utilizing additional features such as the Laplacian positional encoding. In our evaluation of SWAN, we chose not to use additional feature enhancements, in order to provide a clear exposition of the contribution and importance of the local and global non-dissipativity offered by SWAN. The additional comparisons are given in Table 7. In the Peptide-struct task, SWAN demonstrates superior performance compared to all the compared methods. On the Peptide-func task, SWAN outperforms MPNNs, Graph Transformers, DE-GNNs, and some multi-hop GNNs performance to other methods, while being second to multi-hop methods such as DRew. The PascalVOC-SP results show the effectiveness of SWAN with other DE-GNNs (GRAND, GraphCON, ADGN), which are the focus of our paper, while offering competitive results to multi-hop and transformer methods. However, it is essential to note that multi-hop GNNs incur higher complexity, while SWAN maintains a linear complexity. Therefore, we conclude that SWAN offers a highly effective approach for tasks that require long-range interactions, as in the LRGB benchmark.
Table 7: Performance of various classical, multi-hop and static rewiring MPNN, graph Transformer benchmarks, DE-GNNs, and our SWAN across two LRGB tasks. Results are averaged over 3 weight initializations. The first, second, and third best results for each task are color-coded. Beseline results are reported from (Gutteridge et al., 2023).
@@
GCN 0.5930±0.0023 0.3496±0.0013 0.1268±0.0060
GINE 0.5498±0.0079 0.3547±0.0045 0.1265±0.0076
GCNII 0.5543±0.0078 0.3471±0.0010 0.1698±0.0080
GatedGCN 0.5864±0.0077 0.3420±0.0013 0.2873±0.0219
GatedGCN+PE 0.6069±0.0035 0.3357±0.0006 0.2860±0.0085
@@
DIGL+MPNN 0.6469±0.0019 0.3173±0.0007 0.2824±0.0039
DIGL+MPNN+LapPE 0.6830±0.0026 0.2616±0.0018 0.2921±0.0038
MixHop-GCN 0.6592±0.0036 0.2921±0.0023 0.2506±0.0133
MixHop-GCN+LapPE 0.6843±0.0049 0.2614±0.0023 0.2218±0.0174
DRew-GCN 0.6996±0.0076 0.2781±0.0028 0.1848±0.0107
DRew-GCN+LapPE 0.7150±0.0044 0.2536±0.0015 0.1851±0.0092
DRew-GIN 0.6940±0.0074 0.2799±0.0016 0.2719±0.0043
DRew-GIN+LapPE 0.7126±0.0045 0.2606±0.0014 0.2692±0.0059
DRew-GatedGCN 0.6733±0.0094 0.2699±0.0018 0.3214±0.0021
DRew-GatedGCN+LapPE 0.6977±0.0026 0.2539±0.0007 0.3314±0.0024
@@
Transformer+LapPE 0.6326±0.0126 0.2529±0.0016 0.2694±0.0098
SAN+LapPE 0.6384±0.0121 0.2683±0.0043 0.3230±0.0039
GraphGPS+LapPE 0.6535±0.0041 0.2500±0.0005 0.3748±0.0109
@@
GRAND 0.5789±0.0062 0.3418±0.0015 0.1918±0.0097
GraphCON 0.6022±0.0068 0.2778±0.0018 0.2108±0.0091
ADGN 0.5975±0.0044 0.2874±0.0021 0.2349±0.0054
@@
SWAN 0.6313±0.0046 0.2571±0.0018 0.2796±0.0048
SWAN-learn 0.6751±0.0039 0.2485±0.0009 0.3192±0.0250

### v2

@@ D.1 Graph Transfer Dataset
We built the graph transfer datasets upon (Di Giovanni et al. 2023). In each task, graphs use identical topology, but, differently from the original work, nodes are initialized with random input features sampled from a uniform distribution in the interval $[0,0.5)$. In each graph, we selected a source node and target node and initialized them with labels of value “1” and “0”, respectively. We sampled graphs from three graph distributions, i.e., line, ring, and crossed-ring. Figure 4 shows a visual exemplification of the three types of graphs when the distance between the source and target nodes is 5. Specifically, ring graphs are cycles of size $n$, in which the target and source nodes are placed at a distance of $\lfloor n/2\rfloor$ from each other. Crossed-ring graphs are also cycles of size $n$, but include crosses between intermediate nodes. Even in this case, the distance between source and target nodes remains $\lfloor n/2\rfloor$. Lastly, the line graph contains a path of length $n$ between the source and target node. In our experiments, we consider a regression task, whose aim is to swap source and target node labels while maintaining intermediate nodes unchanged. We use an input dimension of 1, and the distance between source and target nodes is equal to 3, 5, 10, and 50. We generated 1000 graphs for training, 100 for validation, and 100 for testing.
@@ D.2 Graph Property Prediction
In our experiments on graph property prediction, we followed the data generation procedure outlined in (Gravina, Bacciu, and Gallicchio 2023). Graphs were randomly selected from various graph distributions, i.e., Erdős–Rényi, Barabasi-Albert, grid, caveman, tree, ladder, line, star, caterpillar, and lobster. A visual examplification of each distribution is provided in Figure 5. Each graph contained between 25 and 35 nodes, where each node is assigned with random identifiers as input features sampled from a uniform distribution in the interval $[0,1)$. The target values represented single-source shortest paths, node eccentricity, and graph diameter. The dataset comprised a total of 7040 graphs, with 5120 used for training, 640 for validation, and 1280 for testing.
(a) Erdős–Rényi
(b) Barabasi-Albert
(c) Grid
(d) Caveman
(e) Tree
(f) Ladder
(g) Line
(h) Star
(i) Caterpillar
(j) Lobster
Figure 5: Erdős–Rényi, Barabasi-Albert, grid, caveman, tree, ladder, line, star, caterpillar, and lobster graphs where the number of nodes is equal to 35.
@@ D.3 Long Range Graph Benchmark
In the Long Range Graph Benchmark section, we considered the “Peptides-func”, “Peptides-struct”, and “PascalVOC-sp” datasets (Dwivedi et al. 2022). In the first two datasets, the graphs correspond to 1D amino acid chains, and they are derived such that the nodes correspond to the heavy (non-hydrogen) atoms of the peptides while the edges represent the bonds between them. Peptides-func is a multi-label graph classification dataset, with a total of 10 classes based on the peptide function, e.g., Antibacterial, Antiviral, cell-cell communication, and others. Peptides-struct is a multi-label graph regression dataset based on the 3D structure of the peptides. Specifically, the task consists of the prediction of the inertia of the molecules using the mass and valence of the atoms, the maximum distance between each atom-pairs, sphericity, and the average distance of all heavy atoms from the plane of best fit. Both Peptides-func and Peptides-struct consist of 15,535 graphs with a total of 2.3 million nodes. PascalVOC-sp is a node classification dataset composed of graphs created from the images in the Pascal VOC 2011 dataset (Everingham et al. 2015). A graph is derived from each image by extracting superpixel nodes using the SLIC algorithm (Achanta et al. 2012) and constructing a rag-boundary graph to interconnect these nodes. Each node represents a region of an image that belongs to a specific class. The dataset consists of 11,355 graphs for a total of 5.4 million nodes. The task involves predicting the semantic segmentation label for each superpixel node across 21 different classes.
We applied stratified splitting to Peptides-func and Peptides-struct to generate balanced train–valid–test dataset splits, using the ratio of 70%–15%–15%, respectively. On PascalVOC-sp, we consider 8,498 graphs for training, 1,428 for validation, and 1,429 test.
@@ Appendix E Experimental Details
Our method is implemented in PyTorch. Our experimental results were obtained using NVIDIA RTX-3090 and A100 GPUs.
@@ E.1 SWAN versions
In our experiments, we explore four variants of our SWAN method, depending on the implementation of $\Phi$ and $\Psi$, as reported in Table 4. These versions can be grouped into two main groups, which differ in the resulting non-dissipative behavior. In the first group, which contains SWAN and SWAN-learn both $\Phi$ and $\Psi$ lead to purely imaginary eigenvalues, since we use symmetric graph shift operators and antisymmetric weight matrices, thus allowing for a node- and graph-wise non-dissipative behavior. The last group, which includes SWAN-ne and SWAN-learn-ne, can deviate from being globally non-dissipative. Indeed, if the weight matrix, ${\bf V}$, is an arbitrary matrix, then the eigenvalues of the Jacobian matrix of the system, $\mathbf{J}(t)$, are contained in a neighborhood of the imaginary axis with radius $r\leq||\mathbf{V}||$ (Bauer-Fike’s theorem (Bauer and Fike 1960)). Although this result does not guarantee that the eigenvalues of the Jacobian are imaginary, in practice, it crucially limits their position, limiting the dynamics of the system on the graph to show, at most, moderate amplification or loss of signals through the graph.
@@ E.3 Graph Property Prediction
We employ the same datasets, hyperparameter space, and experimental setting presented in (Gravina, Bacciu, and Gallicchio 2023). Therefore, we perform hyperparameter tuning via grid search, optimizing the Mean Square Error (MSE), training the models using Adam optimizer for a maximum of 1500 epochs, and early stopping with patience of 100 epochs on the validation error. For each model configuration, we perform 4 training runs with different weight initialization and report the average of the results. We report in Appendix E.5 the grid of hyperparameters exploited for this experiment, while in Appendix D more details about the dataset.
@@ E.4 Long-Range Graph Benchmarks
We employ the same datasets and experimental setting presented in (Dwivedi et al. 2022). Therefore, we perform hyperparameter tuning via grid search, optimizing the Average Precision (AP) in the Peptide-func task, the Mean Absolute Error (MAE) in the Peptide-struct task, and e macro weighted F1 score in PascalVOC-sp, training the models using AdamW optimizer for a maximum of 300 epochs. For each model configuration, we perform 3 training runs with different weight initialization and report the average of the results. Also, we follow the guidelines in (Dwivedi et al. 2022; Gutteridge et al. 2023) and stay within the 500K parameter budget. We report in Appendix E.5 the grid of hyperparameters exploited for this experiment, while in Appendix D more details about the dataset.
@@ E.5 Hyperparameters
Table 5: The grid of hyperparameters employed during model selection for the graph transfer tasks (Transfer), graph property prediction tasks (GraphProp), and Long Range Graph Benchmark (LRGB). We observe that, in each graph transfer task, we use a number of layers that is equal to the distance between the source and target nodes.
@@ Appendix F Complexity and Runtimes
Complexity Analysis. Our SWAN architecture remains within the computational complexity of MPNNs (e.g., Kipf and Welling (2017); Xu et al. (2019)) and other DE-GNNs such as GRAND (Chamberlain et al. 2021b). Specifically, each SWAN layer is linear in the number of nodes $|\cal{V}|$ and edges $|\cal{E}|$, therefore it has a time complexity of $\mathcal{O}(|\cal{V}|+|\cal{E}|)$. Assuming we compute $L$ steps of the ODE defined by SWAN, the overall complexity is $\mathcal{O}\left(L\cdot(|\cal{V}|+|\cal{E}|)\right)$.
Runtimes. We measure the training and inference (i.e., the epoch-time and test-set time) runtime of SWAN on the Peptides-struct dataset and compare it with other baselines. Our runtimes show that SWAN is able to obtain high performance, while retaining a linear complexity runtime that is parallel to other MPNNs and DE-GNNs. Our measurements are reported in Table 6, and presents the training and inference runtimes in seconds, as well as the obtained metric for that task, the mean-absolute-error (MAE), for reference. The runtimes were measured on an NVIDIA RTX-3090 GPU with 24GB of memory. In all time measurements, we use a batch size of 64, 128 feature channels, and 5 layers. For reference, we also provide the reported downstream task performance of each method.
@@
Table 6: Measured training and inference runtimes in seconds, and the obtained MAE of SWAN and other baselines on the Peptides-func dataset.
To further evaluate the scalability of our SWAN, we present in Table 7 results on larger datasets, i.e., ogbn-arxiv (Hu et al. 2020a) (169,343 nodes; 1,166,243 edges), Questions (Platonov et al. 2023) (48,921 nodes; 153,540 edges), Tolokers (Platonov et al. 2023) (11,758 nodes; 519,000 edges). We report test accuracy and training/inference runtimes (ms) measured on an NVIDIA A100 with 40GB memory. Our results show that GraphGPS reaches out-of-memory (OOM) on larger datasets, while SWAN and GCN show similar runtimes, with SWAN outperforming GCN.
ogbn-arxiv Questions Tolokers
Method Training Inference Acc ($\uparrow$) Training Inference Acc ($\uparrow$) Training Inference Acc ($\uparrow$)
(ms) (ms) (ms) (ms) (ms) (ms)
GCN 219.04 107.50 71.74±2.90 86.04 41.22 76.09±1.27 141.98 81.87 83.64±0.67
GPS OOM OOM OOM OOM OOM OOM 1272.15 532.93 83.51±0.93
SWAN 228.45 116.06 72.98±2.60 86.19 41.39 77.94±1.16 142.60 82.09 84.01±0.54
SWAN-learn 255.84 131.48 73.26±2.80 104.71 59.80 78.12±1.28 167.04 97.98 84.69±0.63
Table 7: Measured training and inference runtimes (ms), and the obtained Accuracy of SWAN and other baselines on the ogbn-arxiv, Questions, and Tolokers datasets.
@@ Appendix G Additional Comparisons
Graph Property Prediction. Although we experimented SWAN with weight sharing to maintain consistency with DE-GNN literature (Chamberlain et al. 2021b; Rusch et al. 2022; Eliasof, Haber, and Treister 2021; Gravina, Bacciu, and Gallicchio 2023), a more general version of the framework with layer-dependent weights, i.e., ${\bf W}^{(\ell)}-({\bf W}^{(\ell)})^{\top}$, ${\bf V}^{(\ell)}$, and ${\bf Z}^{(\ell)}$, is possible. We report in Table 8 the comparison of SWAN and SWAN-learn with baseline methods employing both weight sharing and layer dependent weights configurations. Overall, the experiments demonstrate that while the weight-sharing version of SWAN already delivers consistent improvements over state-of-the-art methods, the introducing layer-dependent weights leads to further improvements, significantly outperforming all baselines. Lastly, we note that while there is occasional overlap in standard deviations between SWAN (employing weight sharing) and ADGN, (i) SWAN and SWAN-learn exhibit superior average performance, (ii) they significantly outperform ADGN on real-world long-range graph benchmarks (see Section 4.3), and (iii) the use of layer-dependent weights further amplifies the performance gap compared to ADGN.
Model Diameter SSSP Eccentricity
MPNNs
  GCN 0.7424±0.0466 0.9499±${}_{\pm 9.18\cdot 10^{-5}}$ 0.8468±0.0028
  GAT 0.8221±0.0752 0.6951±0.1499 0.7909±0.0222
  GraphSAGE 0.8645±0.0401 0.2863±0.1843 0.7863±0.0207
  GIN 0.6131±0.0990 -0.5408±0.4193 0.9504±0.0007
  GCNII 0.5287±0.0570 -1.1329±0.0135 0.7640±0.0355
DE-GNNs
  DGC 0.6028±0.0050 -0.1483±0.0231 0.8261±0.0032
  GRAND 0.6715±0.0490 -0.0942±0.3897 0.6602±0.1393
  GraphCON 0.0964±0.0620 -1.3836±0.0092 0.6833±0.0074
  ADGN -0.5188±0.1812 -3.2417±0.0751 0.4296±0.1003
Ours - weight sharing
SWAN -0.5249±0.0155 -3.2370±0.0834 0.4094±0.0764
SWAN-learn -0.5981±0.1145 -3.5425±0.0830 -0.0739±0.2190
Ours - layer dependent weights
SWAN -0.6381±0.0358 -3.9342±0.1993 -0.2706±0.0948
SWAN-learn -0.5905±0.0372 -3.8258±0.0950 -0.2245±0.0840
Table 8: Mean test set $log_{10}(\mathrm{MSE})$ and std averaged over 4 random weight initializations on the Graph Property Prediction tasks. The lower, the better. First and second best results and best baseline for each task are color-coded.
Long Range Graph Benchmark. In Table 2 we are interested in directly comparing the performance of SWAN and its non-dissipative properties with other MPNNs and DE-GNNs, as well as the common approach of using graph transformer to address long-range interaction modeling. We now provide additional comparisons with other methods that utilize multi-hop information, and are therefore more computationally expensive than our SWAN, while also utilizing additional features such as the Laplacian positional encoding. In our evaluation of SWAN, we chose not to use additional feature enhancements, in order to provide a clear exposition of the contribution and importance of the local and global non-dissipativity offered by SWAN. The additional comparisons are given in Table 9. In the Peptide-struct task, SWAN demonstrates superior performance compared to all the compared methods. On the Peptide-func task, SWAN outperforms MPNNs, Graph Transformers, DE-GNNs, and some multi-hop GNNs performance to other methods, while being second to multi-hop methods such as DRew. The PascalVOC-SP results show the effectiveness of SWAN with other DE-GNNs (GRAND, GraphCON, ADGN), which are the focus of our paper, while offering competitive results to multi-hop and transformer methods. However, it is essential to note that multi-hop GNNs incur higher complexity, while SWAN maintains a linear complexity. Therefore, we conclude that SWAN offers a highly effective approach for tasks that require long-range interactions, as in the LRGB benchmark.
@@
  GCN 59.30±0.23 0.3496±0.0013 12.68±0.60
  GINE 54.98±0.79 0.3547±0.0045 12.65±0.76
  GCNII 55.43±0.78 0.3471±0.0010 16.98±0.80
  GatedGCN 58.64±0.77 0.3420±0.0013 28.73±2.19
  GatedGCN+PE 60.69±0.35 0.3357±0.0006 28.60±0.85
@@
  DIGL+MPNN 64.69±0.19 0.3173±0.0007 28.24±0.39
  DIGL+MPNN+LapPE 68.30±0.26 0.2616±0.0018 29.21±0.38
  MixHop-GCN 65.92±0.36 0.2921±0.0023 25.06±1.33
  MixHop-GCN+LapPE 68.43±0.49 0.2614±0.0023 22.18±1.74
  DRew-GCN 69.96±0.76 0.2781±0.0028 18.48±1.07
  DRew-GCN+LapPE 71.50±0.44 0.2536±0.0015 18.51±0.92
  DRew-GIN 69.40±0.74 0.2799±0.0016 27.19±0.43
  DRew-GIN+LapPE 71.26±0.45 0.2606±0.0014 26.92±0.59
  DRew-GatedGCN 67.33±0.94 0.2699±0.0018 32.14±0.21
  DRew-GatedGCN+LapPE 69.77±0.26 0.2539±0.0007 33.14±0.24
@@
  Transformer+LapPE 63.26±1.26 0.2529±0.0016 26.94±0.98
  SAN+LapPE 63.84±1.21 0.2683±0.0043 32.30±0.39
  GraphGPS+LapPE 65.35±0.41 0.2500±0.0005 37.48±1.09
@@
  GRAND 57.89±0.62 0.3418±0.0015 19.18±0.97
  GraphCON 60.22±0.68 0.2778±0.0018 21.08±0.91
  ADGN 59.75±0.44 0.2874±0.0021 23.49±0.54
@@
  SWAN 63.13±0.46 0.2571±0.0018 27.96±0.48
  SWAN-learn 67.51±0.39 0.2485±0.0009 31.92±2.50
Table 9: Performance of various classical, multi-hop and static rewiring MPNN, graph Transformer benchmarks, DE-GNNs, and our SWAN across two LRGB tasks. Results are averaged over 3 weight initializations. The first, second, and third best results for each task are color-coded. Beseline results are reported from (Gutteridge et al. 2023).